Explain what you understand by the apparent magnitude of a star. How is it related
to the brightness of the star? An object A has an apparent magnitude of –5. Another
object B has an apparent magnitude of –10. Calculate the ratio of their brightness
1
Expert's answer
2019-04-26T11:27:35-0400

Apparent magnitude of a star is a measure of its brightness in logarithmic units. By conventional definition, apparent magnitude mm is related to the observed flux density F of light from the star (which is its brightness) by the formula

"m = - 5 \log_{100} \frac{F}{F_0} = - 2.5 \log_{10} \frac{F}{F_0} \, ,"

where "F_0" is the reference flux density corresponding to zero apparent magnitude. Since the flux density (brightness) for a star can be measured in different spectral bands (ultraviolet, visible, infrared etc.), one can speak of the apparent magnitude in the corresponding spectral band. For two objects A and B with apparent magnitudes "m_A" and "m_B", respectively, we have

"m_A - m_B = - 5 \log_{100} \frac{F_A}{F_0} + 5 \log_{100} \frac{F_B}{F_0} = 5 \log_{100} \frac{F_B}{F_A} \, ."

Thus, for "m_A = - 5" and "m_B = - 10", we obtain "5 \log_{100} \left( F_B / F_A \right) = 5", or "\log_{100} \left( F_B / F_A \right) = 1". Hence, the ratio of their brightness is "F_B \/ F_A = 100"